# LLMs generate structurally realistic social networks but overestimate political homophily

### v1

@@ Appendix B Additional results
In Table 5, we report the top interests per demographic group. The LLM generates interests as a comma-separated list (see examples in Table 2), which makes it straightforward to separate the list into individual interests and compute, for each group and interest, what percentage of personas in that group have that interest. For this experiment, we use 1,000 personas, instead to our usual set of 50 personas, so that we can estimate percentages over larger populations. We report percentages for all groups with at least 30 personas (see the counts per group in Table 7). We find that “social justice” is a common interest across groups since it dominates Democrats’ interests (62.5%) and Democrats account for a large portion of every other group besides Republicans. We can also see the result of correlations between demographic identities: for example, men, White, and religious populations are likelier to be Republicans, resulting in higher percentages for “conservative politics”.
Real-world homophily for other demographics.
In the main text, we showed that levels of political homophily predicted by the LLM are unrealistically high, compared to reported levels of homophily in real-world social networks (Table 1). What about levels of homophily for the other demographic variables? Across the six models we test, the LLM consistently predicts homophily to be the strongest for political affiliation, followed by religion, then gender and race/ethnicity (Appendix A). Several works agree that gender homophily should be smaller, compared to other demographic variables. For example, McPherson et al. (2001) says, “By the time that they are adults, people have friendship and confidant networks that are relatively sex-integrated (at least when compared to other dimensions like race, age, and education),” and Thelwall (2009) says, “For U.S. friendship in the last century, the key factors, in decreasing order, seem to be race and ethnicity, age, religion, educational level, occupation, and gender.”
However, these quotes also reveal that homophily for race/ethnicity should not be as low as it is for gender. For example, data from Thelwall (2009), who studies connections on MySpace, suggests same-group ratios of 1.96 for race/ethnicity, 1.69 for religion, and 0.96 for gender. Smith et al. (2014), using data from the US General Social Survey (GSS), finds ratios of 1.63 for religion, 1.32 for race/ethnicity, and 1.19 for gender in 1985 and ratios of 2.08 for religion, 1.47 for race/ethnicity, and 1.12 for gender in 2004. These numbers reveal how much variation there is in levels of homophily across studies and over time, making it difficult to evaluate whether the LLM’s levels of homophily are realistic or not. However, two trends that are clear are that gender homophily is relatively low, which the LLM does capture (although it seems to slightly overestimate), and that religion and race/ethnicity homophily should both be higher than gender homophily, the first of which the LLM captures and the second of which the LLM misses. Furthermore, the LLM’s levels of race/ethnicity homophily are 1.23 for Local and 1.22 for Sequential (using GPT-3.5 Turbo), which are below the three real-world numbers we found (1.96, 1.32, and 1.47), suggesting that the LLM may be underestimating race/ethnicity homophily, in addition to overestimating political homophily.
Model costs and comprehension.
Global, demos $O(ND)$ 607 95.267 1 1.133
Local, demos $O(N^{2}D)$ 30450 196.767 50 50
Sequential (friend list), demos $O(N^{2}(D+E))$ 42601.233 335.633 50 50
Sequential (degree), demos $O(N^{2}(D+N))$ 37735 200.300 50 50.033
Sequential (degree), interests $O(N^{2}(I+N))$ 37035 257.967 50 50.033
Sequential (degree), demos + interests $O(N^{2}(D+I+N))$ 57385 190.733 50 50.067
Table 6: Comparing model costs and comprehension over different LLM methods for social network generation. For Sequential, we consider two versions: listing each persona’s current friend list vs. only their current degree. $N$ is the number of nodes in the network, $E$ is the number of edges, and $D$ and $I$ are the number of tokens needed to describe all demographics and interests, respectively.
In this work, we focus primarily on realism and bias to evaluate the generated networks. However, when using LLMs, measures such as token costs and model comprehension are also very important. We measure token cost as the number of input tokens and output tokens, summed over generating the entire network. We also conduct a big-O analysis of how the input tokens scale with network size, based on $N$ (the number of nodes in the network), $E$ (the number of edges), $D$ (the number of tokens to describe all demographics), and $I$ (the number of tokens to describe all interests). To measure model comprehension, we consider the expected number of turns (prompt and response) to generate the entire network, compared to the actual number of turns. The actual exceeds the expected if we could not parse the model’s response (e.g., it gave an ID that does not exist in the network). So that we can measure the concepts of cost and model comprehension independently, we only add to the number of input/output tokens on successful turns.
We report results in Table 6, with all measures averaged over the 30 networks generated per method. The Global method requires, by far, the fewest number of input tokens, since we only need to list all personas once. However, Global has the highest rate of errors, resulting in an actual number of turns that is 13% higher than the expected number of turns. The Local method and Sequential method scale similarly, since they both require iterating through all $N$ personas, and for each persona, presenting them with information about the other $N-1$ personas. Their main difference is what information is provided about each persona: Sequential provides network information, which can be each person’s friend list (resulting in an additional $E$ per iteration) or only each person’s degree (resulting in an additional $N$ per iteration). Both Local and Sequential have good model comprehension, with the models rarely failing and actual turns being only 0-6.7% above expected turns. Overall, this table lays out tradeoffs, since the Global method requires far fewer tokens, but as we saw in the main text (Figures 2-4), it produces much less realistic networks, and it is more prone to invalid responses.
@@
(a) Comparing structural network metrics.
(b) Comparing homophily.
Figure 13: In our experiments, we use a default temperature of 0.8. Our main results do not change significantly if we use a temperature of 0.6 or 1.0 instead.
(a) Comparing structural network metrics.
(b) Comparing homophily.
Figure 14: Our results do not change siginificantly with minor changes to the prompt. We try adding “Pay attention to all demographics” (+“all”) and prompting the LLM to give a short reason for each friend that it selects (+“reason”).
In our experiments, we use a default temperature of 0.8. In Figure 13, we show that our main results do not significantly change if we use a temperature of 0.6 or 1.0 instead. Since the LLM overemphasizes political homophily, we also try adding “Pay attention to all demographics” to the system prompt. We include GPT-4o in this study as well due to its extreme levels of political homophily (Figure 7). In Figure 14, we show that adding this prompt does not significantly change results for GPT-3.5 Turbo or GPT-4o. Finally, as mentioned before, we also find that results do not significantly change if we prompt the LLM to generate a short reason for each friend that it selects. Thus, our results are robust to these perturbations in temperature and prompt.
@@ C.1 Persona construction
Table 7: Marginal distributions of personas’ demographics. Count in 50 indicates the number of personas belonging to each group in the sample of 50 personas we used for most experiments. Count in 1,000 indicates their number in the sample of 1,000 personas we used for evaluating top interests per demographic group (Table 5)
Figure 15: Age distribution in sample of 1,000 personas.
As described in the main text, we include gender, age, race/ethnicity, religion, and political affiliation. In Table 7, we list the number of personas in each demographic group for the sample of 50 personas we used in most of our experiments, as well as the sample of 1,000 personas we used for evaluating top interests per demographic (Table 5). In Figure 15, we visualize the distribution of ages in the sample of 1,000 personas. Below, we explain how we sampled demographic variables per persona.
@@
First, we use data from the US Census US Census Bureau (2023), who provide monthly population estimates for sex, race/ethnicity, and age (individual years, from 0 to 100 years old). Specifically, we downloaded nc-est2023-alldata-r-file07.csv from US Census datasets archived online [^8] https://www2.census.gov/programs-surveys/popest/datasets/2020-2023/national/asrh/ and used the data for June 2023. We use these estimates to calculate joint distributions of gender, race/ethnicity, and age. Additionally, using data from Pew Research Center Brown (2022), we sample from the age-dependent distribution of those who identify as non-binary.
@@
We sample religion conditioned on the persona’s race/ethnicity. Statista Statista (2016) provides distributions of religious identity for adults in the US in 2016, for most race/ethnicities. Additionally, using data from 2020 PRRI Census of American Religion PRRI Staff (2021), we acquire the distribution for Native Americans.
@@
Finally, we sample political affiliation conditioned on the persona’s race/ethnicity and gender. We primarily use data from Pew Research Center Pew Research Center (2024), using the 2023 numbers from their figure, “Partisan identification by gender among racial and ethnic groups,” which cover most race/ethnicities. Additionally, we use data from Brookings Sanchez and Foxworth (2022), who report Native Americans’ distribution of political support in 2022.
@@
User: In 8-12 words, describe the interests of someone with the following demographics: race/ethnicity: White age: 72 gender: Man political affiliation: Republican religion: Catholic Answer by providing ONLY their interests. Do not include filler like "She enjoys" or "He has a keen interest in".
Figure 16: Prompt to generate interests for persona.
In Figure 16, we provide the prompt that we use to generate interests. We randomize the order of demographics provided, since we find that the LLM seems to pay special attention to the first listed demographic when generating interests. We use GPT-4o to generate interests, since we find that it follows the required format a little better.
@@ C.2 Network generation
Figure 17: Prompt for Global method.
System: You are a Man, age 48, Hispanic, Protestant, Democrat. You are joining a social network. You will be provided a list of people in the network, where each person is described as "ID. Gender, Age, Race/ethnicity, Religion, Political affiliation". Which of these people will you become friends with? Provide a list of *YOUR* friends in the format ID, ID, ID, etc. Do not include any other text in your response. Do not include any people who are not listed below. User: 11. Man, age 31, White, Protestant, Democrat 10. Man, age 58, Hispanic, Catholic, Democrat 41. Woman, age 41, White, Catholic, Republican 2. Woman, age 20, White, Catholic, Republican
Figure 18: Prompt for Local method.
System: You are a Man, age 48, Hispanic, Protestant, Democrat. You are joining a social network. You will be provided a list of people in the network, where each person is described as "ID. Gender, Age, Race/ethnicity, Religion, Political affiliation", followed by their current number of friends. Which of these people will you become friends with? Provide a list of *YOUR* friends in the format ID, ID, ID, etc. Do not include any other text in your response. Do not include any people who are not listed below. User: 11. Man, age 31, White, Protestant, Democrat; has 4 friends 10. Man, age 58, Hispanic, Catholic, Democrat; has 2 friends 41. Woman, age 41, White, Catholic, Republican; has 0 friends 2. Woman, age 20, White, Catholic, Republican; has 7 friends
Figure 19: Prompt for Sequential method, when only degree is provided. We also have a version where each persona’s current list of friends (in the form of ID) is provided.
In Figures 17-19, we provide the full basic prompts for each of our network generation methods: Global, Local, and Sequential. These are the prompts that are used when only demographic variables are provided per persona. When interests are provided, we add “interests include: …” per persona. In variants of the Sequential prompt, discussed in Appendices A and B, we experiment with specifying the number of friends that should be chosen; prompting the LLM to generate a short reason for each selected friend; and adding “Pay attention to all demographics” to the prompt.
@@ C.3 Evaluating network structure
In this section, we provide more details about the real social networks that we compared against and define various measures we used to characterize the networks. In Table 8, we provide basic statistics about the real networks.
@@
Table 8: Summary statistics of the eight real social networks that we use.
@@
We use the following six networks from the CASOS repository [^9] http://www.casos.cs.cmu.edu/tools/datasets/external/index.php:
Galesburg Coleman et al. (1957). This network describes friendship ties between physicians, where they were asked to name three doctors whom they considered personal friends and to nominate three doctors with whom they would discuss medical matters. The goal of this study was to analyze the diffusion of a new drug in terms of when physicians first prescribed it, studied in the context of their social network.
Hi-tech Krackhardt (1999). This network describes friendship ties between employees of a small hi-tech firm. In a survey, they answered the question, “Who do you consider to be a personal friend?” Most friendship nominations were reciprocated, and an edge is kept only if both people nominated each other.
Karate Zachary (1977). This network describes friendships between members of a karate club at a US university. Due to a schism where the club split into two, this network has often been used to study community structure.
Prison MacRae Jr. (1960). This network describes friendship ties between prison inmates. All were asked, “What fellows on the tier are you closest friends with?” Each respondent could choose as many or as few friends as he desired.
Tailor Kapferer (1972). This network describes relations between workers at a tailor shop in Zambia (then Northern Rhodesia). The dataset includes both “instrumental” ties (work-related), which we leave out, and “sociational” ties (friendship, socioemotional), which we include. Networks were recorded twice, seven months apart, so we have two networks from this dataset.
We also include two networks from KONECT [^10] http://konect.uni-koblenz.de/networks/:
@@
$Q=\frac{1}{2E}\sum_{ij}(A_{ij}-\gamma\frac{N_{i}N_{j}}{2E})\mathbbm{1}[c_{i}=c_{j}],$ (3)
@@
We also need to define how we quantified the distance between the generated networks and real networks in Table 3. Let $x_{1},\cdots,x_{m}$ represent the values of a metric (e.g., density) from the real networks (where $m=8$), and let $y_{1},\cdots,y_{n}$ represent the values of the metric from the generated networks (where $n=30)$. First, we compute the difference in their mean, divided by the standard deviation of the real network distribution:
@@
We normalize by standard deviation to make differences comparable across metrics. We only normalize by the real networks’ standard deviation since normalizing by the generated networks’ standard deviation would arbitrarily reward higher variance methods. Second, we use the two-sample Kolmogorov–Smirnov (KS) statistic, which measures the distance between two empirical distributions by comparing their cumulative distribution functions Hodges Jr. (1958):
$F_{\textrm{real}}(u)$ $=\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}[x_{i}\leq u]$
$F_{\textrm{gen}}(u)$ $=\frac{1}{n}\sum_{j=1}^{n}\mathbbm{1}[y_{j}\leq u]$
@@
Erdős–Rényi random graph Erdös and Rényi (1959). We use the $G_{N,p}$ random graph model, which has $N$ nodes and each edge is included with independent probability $p$. In our experiments, we simply set $p$ to the average density of the eight real social networks.
Barabási-Albert preferential attachment Barabási and Albert (1999). In this model, one node is added to the graph at each step, and it forms $m$ edges with existing nodes, where each neighbor is sampled with probability proportional to its current degree. The networkx implementation starts by default with a star graph on $m+1$ nodes, then adds the remaining $N-m-1$ nodes one at a time. Thus, the number of edges in the graph is always $m+((N-m-1)\cdot m)$. We choose $m=5$, which minimizes the difference between the generated graph’s density and the average density of the real social networks.
Watts-Strogatz small world Watts and Strogatz (1998). In this model, first a ring is created over $N$ nodes, then each node is joined to its $k$ nearest neighbors, forming a lattice. Then, with independent probability $p$, each edge $(i,j)$ is rewired, meaning it is replaced with $(i,j^{\prime})$, where $j^{\prime}$ is selected from all nodes (aside from $i$ and $i$’s existing neighbors) uniformly at random. Since each node is joined to its $k$ nearest neighbors, the number of edges in this graph is always $\frac{nk}{2}$. So, we choose $k=10$, which also minimizes the difference between the generated and real networks’ average density. Then, with $N=50$ and $k=10$, we sweep over possible values of $p$ in $\{0.01,0.02,\cdots,0.5\}$ to minimize the difference between the generated and real networks’ average clustering coefficient, resulting in $p=0.15$.

### v2

@@ Appendix B Additional Results
In Table B4, we report the top interests per demographic group. The LLM generates interests as a comma-separated list (see examples in Table 3), which makes it straightforward to separate the list into individual interests and compute, for each group and interest, what percentage of personas in that group have that interest. For this experiment, we use 1,000 personas, instead to our usual set of 50 personas, so that we can estimate percentages over larger populations. We report percentages for all groups with at least 30 personas (see the counts per group in Table C1). We find that “social justice” is a common interest across groups since it dominates Democrats’ interests (62.5%) and Democrats account for a large portion of every other group besides Republicans. We can also see the result of correlations between demographic identities: for example, men, White, and religious populations are likelier to be Republicans, resulting in higher percentages for “conservative politics”.
@@
Figure B5: In our experiments, we use a default temperature of 0.8. Our main results do not change significantly if we use a temperature of 0.6 or 1.0 instead. Top: structural network metrics. Bottom: homophily, where ratios below 1 (marked by the grey line) indicate homophily and lower ratios indicate more homophily. In both plots, we visualize mean and standard error (in black) and individual data points corresponding to each network. All model results shown here use the Sequential method and GPT-3.5-Turbo.
Figure B6: Our results do not change siginificantly with minor changes to the prompt. We try adding “Pay attention to all demographics” (+“all”) and prompting the LLM to give a short reason for each friend that it selects (+“reason”). Top: structural network metrics. Bottom: homophily, where ratios below 1 (marked by the grey line) indicate homophily and lower ratios indicate more homophily. In both plots, we visualize mean and standard error (in black) and individual data points corresponding to each network. All model results shown here use the Sequential method.
In our experiments, we use a default temperature of 0.8. In Figure B5, we show that our main results do not significantly change if we use a temperature of 0.6 or 1.0 instead. Since the LLM overemphasizes political homophily, we also try adding “Pay attention to all demographics” to the system prompt. We include GPT-4o in this study as well due to its extreme levels of political homophily (Figure A1). In Figure B6, we show that adding this prompt does not significantly change results for GPT-3.5 Turbo or GPT-4o. Finally, as mentioned before, we also find that results do not significantly change if we prompt the LLM to generate a short reason for each friend that it selects. Thus, our results are robust to these perturbations in temperature and prompt.
@@ C.1 Persona construction
Table C1: Marginal distributions of personas’ demographics. Count in 50 indicates the number of personas belonging to each group in the sample of 50 personas we used for most experiments. Count in 1,000 indicates their number in the sample of 1,000 personas we used for evaluating top interests per demographic group (Table B4)
Figure C1: Age distribution in sample of 1,000 personas.
As described in the main text, we include gender, age, race/ethnicity, religion, and political affiliation. In Table C1, we list the number of personas in each demographic group for the sample of 50 personas we used in most of our experiments, as well as the sample of 1,000 personas we used for evaluating top interests per demographic (Table B4). In Figure C1, we visualize the distribution of ages in the sample of 1,000 personas. Below, we explain how we sampled demographic variables per persona.
@@
First, we use data from the US Census (US Census Bureau 2023), who provide monthly population estimates for sex, race/ethnicity, and age (individual years, from 0 to 100 years old). Specifically, we downloaded nc-est2023-alldata-r-file07.csv from US Census datasets archived online [^6] https://www2.census.gov/programs-surveys/popest/datasets/2020-2023/national/asrh/ and used the data for June 2023. We use these estimates to calculate joint distributions of gender, race/ethnicity, and age. Additionally, using data from Pew Research Center (Brown 2022), we sample from the age-dependent distribution of those who identify as non-binary.
@@
We sample religion conditioned on the persona’s race/ethnicity. Statista (Statista 2016) provides distributions of religious identity for adults in the US in 2016, for most race/ethnicities. Additionally, using data from 2020 PRRI Census of American Religion (PRRI Staff 2021), we acquire the distribution for Native Americans.
@@
Finally, we sample political affiliation conditioned on the persona’s race/ethnicity and gender. We primarily use data from Pew Research Center (Pew Research Center 2024), using the 2023 numbers from their figure, “Partisan identification by gender among racial and ethnic groups,” which cover most race/ethnicities. Additionally, we use data from Brookings (Sanchez and Foxworth 2022), who report Native Americans’ distribution of political support in 2022.
@@
User: In 8-12 words, describe the interests of someone with the following demographics: race/ethnicity: White age: 72 gender: Man political affiliation: Republican religion: Catholic Answer by providing ONLY their interests. Do not include filler like “She enjoys” or “He has a keen interest in”.
Figure C2: Prompt to generate interests for persona.
In Figure C2, we provide the prompt that we use to generate interests. We randomize the order of demographics provided, since we find that the LLM seems to pay special attention to the first listed demographic when generating interests. We use GPT-4o to generate interests, since we find that it follows the required format a little better.
@@ C.2 Network generation
Figure C3: Prompt for Global method.
System: You are a Man, age 48, Hispanic, Protestant, Democrat. You are joining a social network. You will be provided a list of people in the network, where each person is described as “ID. Gender, Age, Race/ethnicity, Religion, Political affiliation”. Which of these people will you become friends with? Provide a list of *YOUR* friends in the format ID, ID, ID, etc. Do not include any other text in your response. Do not include any people who are not listed below. User: 11. Man, age 31, White, Protestant, Democrat 10. Man, age 58, Hispanic, Catholic, Democrat 41. Woman, age 41, White, Catholic, Republican 2. Woman, age 20, White, Catholic, Republican
Figure C4: Prompt for Local method.
System: You are a Man, age 48, Hispanic, Protestant, Democrat. You are joining a social network. You will be provided a list of people in the network, where each person is described as “ID. Gender, Age, Race/ethnicity, Religion, Political affiliation”, followed by their current number of friends. Which of these people will you become friends with? Provide a list of *YOUR* friends in the format ID, ID, ID, etc. Do not include any other text in your response. Do not include any people who are not listed below. User: 11. Man, age 31, White, Protestant, Democrat; has 4 friends 10. Man, age 58, Hispanic, Catholic, Democrat; has 2 friends 41. Woman, age 41, White, Catholic, Republican; has 0 friends 2. Woman, age 20, White, Catholic, Republican; has 7 friends
Figure C5: Prompt for Sequential method, when only degree is provided. We also have a version where each persona’s current list of friends (in the form of ID) is provided.
In Figures C3-C5, we provide the full basic prompts for each of our network generation methods: Global, Local, and Sequential. These are the prompts that are used when only demographic variables are provided per persona. When interests are provided, we add “interests include: …” per persona. In variants of the Sequential prompt, discussed in Appendices A and B, we experiment with specifying the number of friends that should be chosen; prompting the LLM to generate a short reason for each selected friend; and adding “Pay attention to all demographics” to the prompt.
@@
In Figure 2, we visualize examples of networks generated by each of the three methods. The networks are visualized using networkx, with a spring layout and fixed seed for the visualization. The visualizations primarily reveal differences in the generated networks across methods, but they also reveal differences between networks generated by the same method. Since the visualization method is fixed, these differences should be attributed to variation in actual networks, driven by the randomization in persona ordering (for listing and assignment) and randomness in LLM outputs (using a default temperature of 0.8).
Model costs and comprehension.
Global, demos $O(ND)$ 607 95.267 1 1.133
Local, demos $O(N^{2}D)$ 30450 196.767 50 50
Sequential (friend list), demos $O(N^{2}D+NE)$ 42601.233 335.633 50 50
Sequential (degree), demos $O(N^{2}D)$ 37735 200.300 50 50.033
Sequential (degree), interests $O(N^{2}I)$ 37035 257.967 50 50.033
Sequential (degree), demos + interests $O(N^{2}(D+I))$ 57385 190.733 50 50.067
Table C2: Comparing model costs and comprehension over different LLM methods for social network generation. For Sequential, we consider two versions: listing each persona’s current friend list vs. only their current degree. $N$ is the number of nodes in the network, $E$ is the number of edges, and $D$ and $I$ are the number of tokens needed to describe all demographics and interests, respectively.
In this work, we focus primarily on realism and bias to evaluate the generated networks. However, when using LLMs, measures such as token costs and model comprehension are also very important. We measure token cost as the number of input tokens and output tokens, summed over generating the entire network. We also conduct a big-O analysis of how the input tokens scale with network size, based on $N$ (the number of nodes in the network), $E$ (the number of edges), $D$ (the number of tokens to describe a persona’s demographics), and $I$ (the number of tokens to describe a persona’s interests). To measure model comprehension, we consider the expected number of turns (prompt and response) to generate the entire network, compared to the actual number of turns. The actual exceeds the expected if we could not parse the model’s response (e.g., it gave an ID that does not exist in the network). So that we can measure the concepts of cost and model comprehension independently, we only add to the number of input/output tokens on successful turns.
We report results in Table C2, with all measures averaged over the 30 networks generated per method. The Global method requires, by far, the fewest number of input tokens, since we only need to list all personas once. However, Global has the highest rate of errors, resulting in an actual number of turns that is 13% higher than the expected number of turns. The Local method and Sequential method scale similarly, since they both require iterating through all $N$ personas, and for each persona, presenting them with information about the other $N-1$ personas. Their main difference is what information is provided about each persona: Local only provides $D$ demographics, while Sequential also provides network information, which can be each persona’s friend list (resulting in an additional $E$ per iteration) or only each persona’s degree (resulting in $D+1$ per persona). Both Local and Sequential have good model comprehension, with the models rarely failing and actual turns being only 0-6.7% above expected turns. Overall, this table lays out tradeoffs, since the Global method requires far fewer tokens, but as we saw in the main text (Figures 2-4), it produces much less realistic networks, and it is more prone to invalid responses.
Scalability.
Figure C6: Comparing original generated networks with 50 personas to generated networks with 300 personas, where 30 personas are sampled uniformly at random per query. Top: structural network metrics. Bottom: homophily, where ratios below 1 (marked by the grey line) indicate homophily and lower ratios indicate more homophily. In both plots, we visualize mean and standard error (in black) and individual data points corresponding to each network. All model results shown here use the Sequential method and GPT-3.5-Turbo.
To address these tradeoffs, we propose a simple extension of the Local and Sequential methods to make them more scalable. Previously, when we queried a given persona, we presented it with information about the $N-1$ other personas. Instead, we could provide it a subset of the other personas, with a fixed subset size $k$. As a result, instead of scaling on the order of $O(N^{2}D)$, Local and Sequential would scale on the order of $O(ND)$, which is the same as the Global method.
To demonstrate how this could be done, we implement a simple version of this extension, where the subset is sampled uniformly at random. Then, using the Sequential method, we generate networks with 300 personas, sampling 30 random personas per query. Compared to the smaller 50-node networks we generated before, these larger networks show similar levels of homophily across the five demographic variables, and political homophily remains by far the strongest (Figure C6, bottom). In terms of structural characteristics, the larger networks match the smaller ones (and the real networks) on average shortest paths and modularity, but the larger networks have lower density and clustering (Figure C6, top). This is expected, due to the sampling: previously, each persona would select friends from 100% of all other personas, but now each persona selects friends from only 10% of other personas. Future work could explore methods of non-uniform sampling that can correct for these lower levels of density and clustering, such as using a recommender system-like model (potentially powered by a graph neural network) to choose the subset of other personas that each persona sees.
@@ C.3 Evaluating network structure
In this section, we provide more details about the real social networks that we compared against and define various measures we used to characterize the networks. In Table C3, we provide basic statistics about the real networks.
@@
Table C3: Summary statistics of the eight real social networks that we use.
@@
We use the following six networks from the CASOS repository (CASOS 2024):
Galesburg (Coleman, Katz, and Menzel 1957). This network describes friendship ties between physicians, where they were asked to name three doctors whom they considered personal friends and to nominate three doctors with whom they would discuss medical matters. The goal of this study was to analyze the diffusion of a new drug in terms of when physicians first prescribed it, studied in the context of their social network.
Hi-tech (Krackhardt 1999). This network describes friendship ties between employees of a small hi-tech firm. In a survey, they answered the question, “Who do you consider to be a personal friend?” Most friendship nominations were reciprocated, and an edge is kept only if both people nominated each other.
Karate (Zachary 1977). This network describes friendships between members of a karate club at a US university. Due to a schism where the club split into two, this network has often been used to study community structure.
Prison (MacRae Jr. 1960). This network describes friendship ties between prison inmates. All were asked, “What fellows on the tier are you closest friends with?” Each respondent could choose as many or as few friends as he desired.
Tailor (Kapferer 1972). This network describes relations between workers at a tailor shop in Zambia (then Northern Rhodesia). The dataset includes both “instrumental” ties (work-related), which we leave out, and “sociational” ties (friendship, socioemotional), which we include. Networks were recorded twice, seven months apart, so we have two networks from this dataset.
We also include two networks from the KONECT repository (KONECT 2024):
@@
$Q=\frac{1}{2E}\sum_{ij}(A_{ij}-\gamma\frac{N_{i}N_{j}}{2E})\mathds{1}[c_{i}=c_{j}],$ (5)
@@
We also need to define how we quantified the distance between the generated networks and real networks in Table B1. Let $x_{1},\cdots,x_{m}$ represent the values of a metric (e.g., density) from the real networks (where $m=8$), and let $y_{1},\cdots,y_{n}$ represent the values of the metric from the generated networks (where $n=30)$. First, we compute the difference in their mean, divided by the standard deviation of the real network distribution:
@@
We normalize by standard deviation to make differences comparable across metrics. We only normalize by the real networks’ standard deviation since normalizing by the generated networks’ standard deviation would arbitrarily reward higher variance methods. Second, we use the two-sample Kolmogorov–Smirnov (KS) statistic, which measures the distance between two empirical distributions by comparing their cumulative distribution functions (Hodges Jr. 1958):
$F_{\textrm{real}}(u)$ $=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[x_{i}\leq u]$
$F_{\textrm{gen}}(u)$ $=\frac{1}{n}\sum_{j=1}^{n}\mathds{1}[y_{j}\leq u]$
@@
Erdős–Rényi random graph (Erdős and Rényi 1959). We use the $G_{N,p}$ random graph model, which has $N$ nodes and each edge is included with independent probability $p$. In our experiments, we simply set $p$ to the average density of the eight real social networks.
Barabási-Albert preferential attachment (Barabási and Albert 1999). In this model, one node is added to the graph at each step, and it forms $m$ edges with existing nodes, where each neighbor is sampled with probability proportional to its current degree. The networkx implementation starts by default with a star graph on $m+1$ nodes, then adds the remaining $N-m-1$ nodes one at a time. Thus, the number of edges in the graph is always $m+((N-m-1)\cdot m)$. We choose $m=5$, which minimizes the difference between the generated graph’s density and the average density of the real social networks.
Watts-Strogatz small world (Watts and Strogatz 1998). In this model, first a ring is created over $N$ nodes, then each node is joined to its $k$ nearest neighbors, forming a lattice. Then, with independent probability $p$, each edge $(i,j)$ is rewired, meaning it is replaced with $(i,j^{\prime})$, where $j^{\prime}$ is selected from all nodes (aside from $i$ and $i$’s existing neighbors) uniformly at random. Since each node is joined to its $k$ nearest neighbors, the number of edges in this graph is always $\frac{nk}{2}$. So, we choose $k=10$, which also minimizes the difference between the generated and real networks’ average density. Then, with $N=50$ and $k=10$, we sweep over possible values of $p$ in $\{0.01,0.02,\cdots,0.5\}$ to minimize the difference between the generated and real networks’ average clustering coefficient, resulting in $p=0.15$.